Problem 3-3A Preparing adjusting entries, adjusted trial balance, and financial statements LO P1, P2, P3

[The following information applies to the questions displayed below.]

Wells Technical Institute (WTI), a school owned by Tristana Wells, provides training to individuals who pay tuition directly to the school. WTI also offers training to groups in off-site locations. Its unadjusted trial balance as of December 31, 2017, follows. WTI initially records prepaid expenses and unearned revenues in balance sheet accounts. Descriptions of items a through h that require adjusting entries on December 31, 2017, follow.
  
Additional Information Items

  1. An analysis of WTI's insurance policies shows that $3,864 of coverage has expired.
  2. An inventory count shows that teaching supplies costing $3,349 are available at year-end 2017.
  3. Annual depreciation on the equipment is $15,458.
  4. Annual depreciation on the professional library is $7,729.
  5. On November 1, WTI agreed to do a special six-month course (starting immediately) for a client. The contract calls for a monthly fee of $2,900, and the client paid the first five months' fees in advance. When the cash was received, the Unearned Training Fees account was credited. The fee for the sixth month will be recorded when it is collected in 2018.
  6. On October 15, WTI agreed to teach a four-month class (beginning immediately) for an individual for $4,700 tuition per month payable at the end of the class. The class started on October 15, but no payment has yet been received. (WTI's accruals are applied to the nearest half-month; for example, October recognizes one-half month accrual.)
  7. WTI's two employees are paid weekly. As of the end of the year, two days' salaries have accrued at the rate of $100 per day for each employee.
  8. The balance in the Prepaid Rent account represents rent for December.
WELLS TECHNICAL INSTITUTE
Unadjusted Trial Balance
December 31, 2017
Debit Credit
Cash $ 27,245
Accounts receivable 0
Teaching supplies 10,478
Prepaid insurance 15,719
Prepaid rent 2,097
Professional library 31,436
Accumulated depreciation—Professional library $ 9,432
Equipment 73,338
Accumulated depreciation—Equipment 16,768
Accounts payable 34,749
Salaries payable 0
Unearned training fees 14,500
Common stock 13,000
Retained earnings 53,646
Dividends 41,916
Tuition fees earned 106,885
Training fees earned 39,820
Depreciation expense—Professional library 0
Depreciation expense—Equipment 0
Salaries expense 50,300
Insurance expense 0
Rent expense 23,067
Teaching supplies expense 0
Advertising expense 7,336
Utilities expense 5,868
Totals $ 288,800 $ 288,800

Problem 3-3A Part 1

Required:
1. Prepare the necessary adjusting journal entries for items a through h. Assume that adjusting entries are made only at year-end.
Answer #1
Wells Technical Institute WTI
Journal entries
No. Account Debit Credit Calculation
a Insurance expense $        3,864
Prepaid Insurance $        3,864
b Teaching Supplies expense $        7,129 =10,478-3,349  
Teaching Supplies $        7,129
c Depreciation expense- Equipment $     15,458
Accumulated Depreciation-Equipment $     15,458
d Depreciation expense- Prof Library $        7,729
Accumulated Depreciation-Professional Library $        7,729
e Unearned Training Fees $        5,800 =2,900*2
Training fees eanred $        5,800
f Accounts Receivable $     11,750 =4,700*2.5
Tuition fees eanred $     11,750
g Salaries expense $           400 =100*2*2
Salaries Payable $           400
h Rent expense $        2,097
Prepaid Rent $        2,097